# Insulator phases of Bose–Fermi mixtures induced by intraspecies next-neighbor interactions

### v2

@@ Abstract
We study a one-dimensional mixture of two-color fermions and scalar bosons at the hardcore limit, focusing on the effect that the intraspecies next-neighbor interactions have on the zero-temperature ground state of the system for different fillings of each carrier. Exploring the problem’s parameters, we observed that the non-local interaction could favor or harm the well-known mixed Mott and spin-selective Mott insulators. We also found the emergence of three unusual insulating states with charge density wave (CDW) structures in which the orders of the carriers are out of phase with each other. For instance, the immiscible CDW appears only at half-filling bosonic density, whereas the mixed CDW state is characterized by equal densities of bosons and fermions. Finally, the spin-selective CDW couples the bosons and only one kind of fermions. Appropriate order parameters were proposed for each phase to obtain the critical parameters for the corresponding superfluid-insulator transition. Our results can inspire or contribute to understanding experiments in cold-atom setups with long-range interactions or recent reports involving quasiparticles in semiconductor heterostructures.
@@ 1 Introduction
The development of quantum simulators over the last few decades has allowed the control of many-body quantum systems’ interactions and dynamics to a great extent [1, 2]. This has paved the way for understanding the physics behind Bose–Fermi mixtures: composition of particles with both bosonic and fermionic statistics, which have been studied for decades, with one of the first examples being the ${}^{3}\rm{He}$-${}^{4}\rm{He}$ combination [3, 4, 5]. Here the development of cold-atom simulators has been crucial [6, 7, 8], as it has been possible to produce an amalgam of degenerate mixtures of gases [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] that exhibit novel and fascinating phenomena such as Bose–Fermi superfluidity [27, 28, 29, 30] or dual Mott insulators [31]. Even so, recent experiments of dipolar excitons in 2D materials [32] present a promising alternative in the simulation of long-range interacting Bose–Fermi mixtures, since these quasiparticles possess an inherent non-local interaction that enriches the physics of the system [33, 34, 35, 36, 37].
@@
To perform a theoretical study of these mixtures, it is usual to describe fermions and bosons using the Hubbard model and coupling terms that account for interspecies interactions. Within this framework, there have been abundant numerical and analytical studies, which have resulted in the prediction of diverse kinds of carrier configurations, such as Luttinger liquids, charge density waves (CDW), and Mott insulators (MI), among others [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. Of particular interest for this research are the superfluid-insulator transitions, where it is well known that the insulator phases are characterized by commensurable relations between the fixed fermionic $\left(\rho_{\rm{F}}\right)$ and bosonic $\left(\rho_{\rm{B}}\right)$ densities. Specifically, for a two-color fermion and scalar boson mixture with repulsive interspecies interactions, there is always one mixed Mott insulator (MMI) characterized by the relation $\rho_{\rm{B}}+\rho_{\rm{F}}=n$ ($n$ being an integer) and a spin-selective Mott insulator (SSMI) that follows $\rho_{\rm{B}}+\rho^{\uparrow,(\downarrow)}_{\rm{F}}=n$, which indicates that for a population imbalance, this last insulator is divided in two [64, 65, 66, 67, 68]. It is important to note that flavor-selective Mott phases have been observed experimentally in $\rm{SU}(3)$ Fermi Hubbard models where the symmetry is explicitly broken by adding an inter-band hopping [69].
@@
Non-local interactions have enriched the ground-state phase diagram of both bosonic and fermionic systems, allowing diverse carrier configurations. In particular, adding next-neighbor interactions to a bosonic system has led to the emergence of exotic phases, including supersolid, Haldane, and solitonic ones, and the charge density wave phase for integer and half-integer densities [70, 71, 72, 73, 74]. On the other hand, the one-dimensional phase diagram of fermions under local and non-local interactions is well known for half-filling. However, it is controversial and little studied for other fillings [75, 76, 77, 78, 79, 80, 81, 82]. At half-filling, a charge density wave phase emerges when the long-range interaction prevails over the on-site one, and a spin density wave appears when the opposite happens. In addition, other phases appear, such as a bond density wave when both interactions are of the same order, a phase separation for attractive interactions, and certain superconductor phases, singlet and triplet.
@@
Despite the interesting physics generated by the non-local interactions, there are only a few studies where this effect is considered for a Bose–Fermi mixture, with one example being an investigation with polarized fermions [83]. Hence we proposed exploring the extended Bose–Fermi Hubbard model with two-color fermions and scalar bosons at the hardcore limit, including next-neighbor intraspecies interactions. As a result, we obtained a phase diagram for each type of non-local interaction, bosonic or fermionic, which summarize the different insulator phases that can appear at a given filling (see Fig. 1). The latter contains the well-known mixed and spin-selective Mott insulators, as well as three unusual incommensurable insulators that appear when taking into account the next-neighbor intraspecies interactions, all of which have charge density wave (CDW) orders for each species that are out of phase with each other. They correspond to the immiscible charge density wave (ICDW), a CDW phase with a period of two sites that only appears for half filling of the bosons $\rho_{\rm{B}}=1/2$, the mixed charge density wave (MCDW) in which both bosons and fermions have the same density, that is $\rho_{\rm{B}}-\rho_{\rm{F}}=0$, and the CDW order is characterized by a wave vector proportional to said density, and the spin-selective charge density wave (SSCDW), which corresponds to a phase analogous to the latter, where the bosons only couple to one of the fermionic spin components $\left(\rho_{\rm{B}}-\rho^{\uparrow,(\downarrow)}_{\rm{F}}=0\right)$.
@@
Figure 1: Phase diagram of the bosonic density $\rho_{\rm{B}}$ vs. the fermionic density $\rho_{\rm{F}}$ for a balanced $(\rho^{\uparrow}_{\rm{F}}=\rho^{\downarrow}_{\rm{F}})$ Bose–Fermi mixture with intraspecies next-neighbor interactions between only fermions (a) and bosons (b), respectively. Here we consider only rational values for both densities, guaranteeing that each density is always commensurate with the chain. In the diagram, the white regions correspond to superfluid phases, while the non-vertical lines indicate the presence of an insulator state, in which we can find the well-known mixed Mott insulator (MMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ (dark blue line) and the spin-selective Mott insulator (SSMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ (dark green line), as well as the three phases found, the immiscible charge density wave (ICDW) at $\rho_{\rm{B}}=1/2$ (purple line), the mixed charge density wave (MCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}=0$ (brown line) and the spin-selective charge density wave (SSCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}/2=0$ (light green line). The vertical lines signal special densities that are related to Fig. 2, 4, 6 and 11. Here, each time these vertical lines intersect with another one an insulator phase appears in the corresponding phase diagram (for suitable values of the couplings). The graph only includes part of the range $\rho_{\rm{F}}>1$, since this region corresponds to a reflection of the $\rho_{\rm{F}}<1$ diagram thanks to the particle-hole symmetry of the Hamiltonian shown in Appendix A.
@@
Therefore, the main objective of this article is to study the conditions under which these incommensurable insulators appear, as well as to expose the main properties that characterize their structure. To achieve this, in Sec. 2 we first present the extended Bose–Fermi Hubbard model to be used, along with a discussion of our numerical techniques, which are matrix product states (MPS) optimization algorithms for finding the ground state. The main results are found in Sec. 3, where we perform an analysis of Fig. 1(a) by choosing three representative fermionic densities, given by $\rho_{\rm F}=2/5$, $2/3$, and $4/5$, each of which shows a particular incommensurable insulator for nonzero fermionic next-neighbor interactions, along with the principal characteristics from the phase diagram. Accordingly, in each subsection, we study the corresponding $\rho_{\rm B}-\mu_{\rm B}$ and $\mu_{\rm B}-V_{\rm FF}$ diagrams in conjunction with appropriate density profiles as a means of analyzing the properties of the different incompressible phases and their dependence on the fermionic next-neighbor interaction, with a particular focus on the incommensurable ones. In Sec. 4, we take a look at Fig. 1(b) and the role of bosonic next-neighbor interactions. Finally, in Sec. 5 we make some final remarks and outline future perspectives for research.
@@ 2 Model
[…]
which is defined on a one-dimensional chain of length $L$ with open boundary conditions for numerical efficiency. In the above Hamiltonian, $\hat{b}^{\dagger}_{j},\hat{b}_{j}$ $\left(\hat{f}^{\dagger}_{\sigma,j},\hat{f}_{\sigma,j}\right)$ are the creator and annihilator boson (fermion with spin $\sigma=\uparrow$, $\downarrow$) operators at the j-th site with their corresponding number operators $\hat{n}^{\rm B}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$, $\hat{n}^{\rm F}_{\sigma,j}=\hat{f}^{\dagger}_{\sigma,j}\hat{f}_{\sigma,j}$, and $\hat{n}^{\rm F}_{j}=\hat{n}^{\rm F}_{\uparrow,j}+\hat{n}^{\rm F}_{\downarrow,j}$. The hopping integral for bosons (fermions) is $t_{\rm B}$ $\left(t_{\rm F}\right)$, where $\langle j,l\rangle$ indicates next-neighbor pairs and $U_{\rm BB}$ $\left(U_{\rm FF}\right)$ is the intensity of the local boson-boson (fermion-fermion) interaction, while the magnitude of the local interspecies interaction is given by $U_{\rm BF}$. Knowing that next-neighbor interactions are relevant, we consider the corresponding intraspecies terms for both fermions and bosons with strengths quantified by $V_{\rm FF}$ and $V_{\rm BB}$, respectively.
Additionally, the previous Hamiltonian admits a set of appropriate Abelian quantum numbers given by the boson number $N_{\rm B}$ and the fermion number $N^{\sigma}_{\rm F}$ with spin $\sigma=\uparrow$, $\downarrow$. From them we define useful quantities for the rest of the study, such as the total fermion number $N_{\rm F}=N_{\rm F}^{\uparrow}+N_{\rm F}^{\downarrow}$, the global boson density $\rho_{\rm B}=N_{\rm B}/L$, and the global fermion densities $\rho_{\rm F}^{\sigma}=N^{\sigma}_{\rm F}/L$, and $\rho_{\rm F}=N_{\rm F}/L$. We also define $I=|\rho_{\rm F}^{\uparrow}-\rho_{\rm F}^{\downarrow}|/\rho_{\rm F}$ as the imbalance of spin population, and then refer to the case of $I=0$ as a balanced mixture.
[…]
To achieve insight into the physics of this model, we perform a ground-state energy search for a given set of particle fillings $E(N_{\rm B},N_{\rm F}^{\uparrow},N_{\rm F}^{\downarrow})$ and different interaction strengths, using a density matrix renormalization group (DMRG) algorithm based on the MPS language [86] with the help of the TeNPy library [87]. We set a maximum truncation error of $10^{-14}$, which varies the bond dimension between sweeps and sites of the chain. On the other hand, since the addition of next-neighbor interactions gives rise to a need for extra degrees of entanglement, we employ a sweep-dependent bond dimension that starts at a given value between 500 and 1,000 and increases each 10 sweeps by 100. The optimization code stops when the energy (entropy) error goes beyond the $10^{-5}$ $\left(10^{-3}\right)$ limit.

### v3

@@ Abstract
We study a one-dimensional mixture of two-color fermions and scalar bosons at the hardcore limit, focusing on the effect that the intraspecies next-neighbor interactions have on the zero-temperature ground state of the system for different fillings of each carrier. Exploring the problem’s parameters, we observed that the nearest-neighbor interaction could favor or harm the well-known mixed Mott and spin-selective Mott insulators. We also found the emergence of three unusual insulating states with charge density wave (CDW) structures in which the orders of the carriers are out of phase between each other. For instance, the immiscible CDW appears only at half-filling bosonic density, whereas the mixed CDW state is characterized by equal densities of bosons and fermions. Finally, the spin-selective CDW couples the bosons and only one kind of fermions. Appropriate order parameters were proposed for each phase to obtain the critical parameters for the corresponding superfluid-insulator transition. Our results can inspire or contribute to understanding experiments in cold-atom setups with long-range interactions or recent reports involving quasiparticles in semiconductor heterostructures.
Copyright attribution to authors.
This work is a submission to SciPost Physics Core.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
The development of quantum simulators over the last few decades has allowed the control of many-body quantum systems’ interactions and dynamics to a great extent [1, 2]. This paved the way to understanding the physics behind Bose–Fermi mixtures: composition of particles with both bosonic and fermionic statistics, which have been studied for decades with one of the first examples being the ${}^{3}\rm{He}$-${}^{4}\rm{He}$ combination [3, 4, 5]. Here the development of cold-atom simulators has been crucial [6, 7, 8], as it has been possible to produce an amalgam of degenerate mixtures of gases [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26] which exhibit novel and fascinating phenomena such as Bose–Fermi superfluidity [27, 28, 29, 30] or dual Mott insulators [31]. Likewise, recent experiments of dipolar excitons in 2D materials [32] present a promising alternative in the simulation of long-range interacting Bose–Fermi mixtures, since this type of interaction is inherent to these quasiparticles which enriches the physics of the system [33, 34, 35, 36, 37, 38, 39].
@@
To perform a theoretical study of these mixtures, it is usual to describe fermions and bosons using the Hubbard model and coupling terms that account for interspecies interactions. Within this framework, there have been abundant numerical and analytical studies, which have resulted in the prediction of diverse kinds of carriers’ configurations, such as Luttinger liquids, charge density waves (CDW), and Mott insulators (MI), among others [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65]. Of particular interest for this research are the superfluid-insulator transitions, where it is well known that the insulator phases are characterized by commensurable relations between the fixed fermionic $\left(\rho_{\rm{F}}\right)$ and bosonic $\left(\rho_{\rm{B}}\right)$ densities. Specifically, for a two-color fermion and scalar boson mixture with repulsive interspecies interactions, there is always one mixed Mott insulator (MMI) characterized by the relation $\rho_{\rm{B}}+\rho_{\rm{F}}=n$ ($n$ being an integer) and a spin-selective Mott insulator (SSMI) that follows $\rho_{\rm{B}}+\rho^{\uparrow,(\downarrow)}_{\rm{F}}=n$, which indicates that for a population imbalance, this last insulator is divided in two [66, 67, 68, 69, 70]. It is important to note that flavor-selective Mott phases have been observed experimentally in $\rm{SU}(3)$ Fermi Hubbard models where the symmetry is explicitly broken by adding an inter-band hopping [71].
@@
Long-ranged interactions have enriched the ground-state phase diagram of both bosonic and fermionic systems allowing diverse carrier configurations. Specifically, adding next-neighbor interactions to a bosonic system led to the emergence of exotic phases including supersolid, Haldane, and solitonic ones, and the charge density wave phase for integer and half-integer densities [72, 73, 74, 75, 76, 77, 78]. On the other hand, the one-dimensional phase diagram of fermions under short and long-ranged interactions is well-known for half-filling. However, it is controversial and little studied for other fillings [79, 80, 81, 82, 83, 84, 85, 86]. At half-filling, a charge density wave phase emerges when the long-range interaction prevails over the on-site one, and a spin density wave when the opposite happens. In addition, other phases appear, such as a bond density wave when both interactions are of the same order, a phase separation for attractive interactions, and certain superconductor phases, singlet and triplet.
Motivated by the latter, we propose an exploration of the extended Bose–Fermi Hubbard model with two-color fermions and scalar bosons at the hardcore limit, including next-neighbor intraspecies interactions. Here we wish to know what new phenomena can emerge from the interplay between long-range interactions and disparity of particle statistics. Moreover, the inclusion of spinful fermions allows the appearance of spin-selective phases, which are also of high interest [66, 67, 68, 69, 70], and the restriction in the form of the hardcore limit keeps a tractable Hilbert space for simulation while also allowing the extrapolation of its results to soft-core models, as shown in previous works [67, 68]. There are only a few studies where these effects are considered for Bose–Fermi mixtures, with one example being an investigation with polarized fermions [87], then we believe this work significantly expands our understanding of this topic. Besides the convergence of on-site and long-range interactions leads us to a multi-parameter model that could be essential to describe physics of recent Bose-Fermi mixtures present in semiconductors heterostructures, where stripes and diverse bidimensional patterns were observed  [33, 34, 35, 36, 37, 38, 39].
@@
In the following, we obtained a phase diagram for each type of long-ranged interaction, bosonic or fermionic, which summarizes the different insulator phases that can appear at a given filling (see Fig. 1). The latter contains the well-known mixed and spin-selective Mott insulators, as well as three unusual insulators that appear when taking into account the next-neighbor intraspecies interactions, all of which have charge density wave (CDW) orders for each species that are out of phase between each other. These correspond to the immiscible charge density wave (ICDW), a CDW phase with a period of two sites that only appears for half filling of the bosons $\rho_{\rm{B}}=1/2$, the mixed charge density wave (MCDW) in which both bosons and fermions have the same density, that is $\rho_{\rm{B}}-\rho_{\rm{F}}=0$, and the CDW order is characterized by a wave vector proportional to said density, and the spin-selective charge density wave (SSCDW) that corresponds to an analogous phase to the latter where the bosons only couple to one of the fermionic spin components $\left(\rho_{\rm{B}}-\rho^{\uparrow,(\downarrow)}_{\rm{F}}=0\right)$.
@@
Figure 1: Phase diagram as a function of the bosonic $\rho_{\rm{B}}$ and fermionic $\rho_{\rm{F}}$ densities for a balanced $(\rho^{\uparrow}_{\rm{F}}=\rho^{\downarrow}_{\rm{F}})$ Bose–Fermi mixture with intraspecies next-neighbor interactions between only fermions (a) and bosons (b), respectively. Here we consider only rational values for both densities, guaranteeing that the number of carriers is always commensurate with the chain. Each non-vertical line corresponds to the density combinations necessary for the emergence of a particular insulator state, while the rest of the diagram are superfluid phases. The insulators correspond to the well-known mixed Mott insulator (MMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ (dark blue line) and the spin-selective Mott insulator (SSMI) at $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ (dark green line), as well as the three phases found: the immiscible charge density wave (ICDW) at $\rho_{\rm{B}}=1/2$ (purple line), the mixed charge density wave (MCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}=0$ (brown line) and the spin-selective charge density wave (SSCDW) at $\rho_{\rm{B}}-\rho_{\rm{F}}/2=0$ (light green line). The vertical lines signal special densities that are related to Fig. 2, 4, 6 and 11, here each time these vertical lines intersect with another non-vertical one an insulator phase appears in the corresponding phase diagram (for suitable values of the couplings). The graph only includes part of the range $\rho_{\rm{F}}>1$, since this region corresponds to a reflection of the $\rho_{\rm{F}}<1$ diagram thanks to the particle-hole symmetry of the Hamiltonian shown in Appendix A.
@@
Therefore, the main objective of this article is to study the conditions under which these CDW insulators appear, as well as to expose the main properties that characterize their structure. To achieve this, in Sec. 2 we first present the extended Bose–Fermi Hubbard model to be used, along with a discussion of our numerical techniques, which are matrix product states (MPS) optimization algorithms for finding the ground state. The main results are found in Sec. 3, where we perform the analysis of Fig. 1(a) by choosing three representative fermionic densities given by $\rho_{\rm F}=2/5$, $2/3$, $4/5$, each of which shows a particular CDW insulator for nonzero fermionic next-neighbor interactions along with the principal characteristics from the phase diagram. Accordingly, in each subsection, we study the corresponding $\rho_{\rm B}-\mu_{\rm B}$ and $\mu_{\rm B}-V_{\rm FF}$ diagrams in conjunction with appropriate density profiles as a means of analyzing the properties of the different incompressible phases and their dependence on the fermionic next-neighbor interaction, with a particular focus on the ones of CDW type. In Sec. 4 we take a look at Fig. 1(b) and the role of bosonic next-neighbor interactions. Finally, in Sec. 5 we give some final remarks and future perspectives for research.
@@ 2 Model
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}_{\rm B}+\hat{\mathcal{H}}_{\rm F}+\hat{\mathcal{H}}_{\rm BF},$ (1)
^$\hat{\mathcal{H}}_{\rm B}=-t_{\rm B}\sum_{\langle j,l\rangle}\left(\hat{b}^{\dagger}_{j}\hat{b}_{l}+\rm{H.c.}\right)+\frac{U_{\rm BB}}{2}\sum_{j}\hat{n}^{\rm B}_{j}\left(\hat{n}^{\rm B}_{j}-1\right)+V_{\rm BB}\sum_{\langle j,l\rangle}\hat{n}^{\rm B}_{j}\hat{n}^{\rm B}_{l},$ (2)
^$\hat{\mathcal{H}}_{\rm F}=-t_{\rm F}\sum_{\langle j,l\rangle,\sigma}\left(\hat{f}^{\dagger}_{\sigma,j}\hat{f}_{\sigma,l}+\rm{H.c.}\right)+U_{\rm FF}\sum_{j}\hat{n}^{\rm F}_{\uparrow,j}\hat{n}^{\rm F}_{\downarrow,j}+V_{\rm FF}\sum_{\langle j,l\rangle}\hat{n}^{\rm F}_{j}\hat{n}^{\rm F}_{l},$ (3)
^$\hat{\mathcal{H}}_{\rm BF}=U_{\rm BF}\sum_{j}\hat{n}^{\rm B}_{j}\hat{n}^{\rm F}_{j},$ (4)
which is defined on a one-dimensional chain of length $L$ with open boundary conditions for numerical efficiency. In the above Hamiltonian $\hat{b}^{\dagger}_{j},\hat{b}_{j}$ $\left(\hat{f}^{\dagger}_{\sigma,j},\hat{f}_{\sigma,j}\right)$ are the creator and annihilator boson (fermion with spin $\sigma=\uparrow$, $\downarrow$) operators at the j-th site with their corresponding number operators $\hat{n}^{\rm B}_{j}=\hat{b}^{\dagger}_{j}\hat{b}_{j}$, $\hat{n}^{\rm F}_{\sigma,j}=\hat{f}^{\dagger}_{\sigma,j}\hat{f}_{\sigma,j}$ and $\hat{n}^{\rm F}_{j}=\hat{n}^{\rm F}_{\uparrow,j}+\hat{n}^{\rm F}_{\downarrow,j}$. The hopping integral for bosons (fermions) is $t_{\rm B}$ $\left(t_{\rm F}\right)$, where $\langle j,l\rangle$ indicates next-neighbor pairs, and $U_{\rm BB}$ $\left(U_{\rm FF}\right)$ is the intensity of the local boson-boson (fermion-fermion) interaction, while the magnitude of the local interspecies interaction is given by $U_{\rm BF}$. Knowing that next-neighbor interactions are relevant, we consider the corresponding intraspecies terms for both fermions and bosons with strengths quantified by $V_{\rm FF}$ and $V_{\rm BB}$, respectively.
Additionally, the previous Hamiltonian admits a set of appropriate Abelian quantum numbers given by the boson number $N_{\rm B}$ and the fermion number $N^{\sigma}_{\rm F}$ with spin $\sigma=\uparrow$, $\downarrow$. From them we define useful quantities for the rest of the study, such as the total fermion number $N_{\rm F}=N_{\rm F}^{\uparrow}+N_{\rm F}^{\downarrow}$, the global boson density $\rho_{\rm B}=N_{\rm B}/L$, and the global fermion densities $\rho_{\rm F}^{\sigma}=N^{\sigma}_{\rm F}/L$, and $\rho_{\rm F}=N_{\rm F}/L$. We also define $I=|\rho_{\rm F}^{\uparrow}-\rho_{\rm F}^{\downarrow}|/\rho_{\rm F}$ as the imbalance of spin population, then we refer to the case of $I=0$ as a balanced mixture.
@@
To achieve insight into the physics of this model, we perform a ground-state energy search for a given set of particle fillings $E(N_{\rm B},N_{\rm F}^{\uparrow},N_{\rm F}^{\downarrow})$ and different interaction strengths, using a two-site density matrix renormalization group (DMRG) algorithm based on the MPS language [90] with the help of the TeNPy library [91]. We employ a sweep-dependent maximum bond dimension which starts at a given value between 500 and 1000 and increases every 10 sweeps by 100. The two-site DMRG algorithm truncates the Schmidt coefficients up to a maximum error of $10^{-14}$ or to the maximum bond dimension of the corresponding sweep, nevertheless, since the addition of next-neighbor interactions gives rise to a need for extra degrees of entanglement the latter situation is the general case for our simulations. The stop criteria for the algorithm are given by the energy and entropy errors given by
$\Delta E_{n}$ $=\left|\frac{E_{n}-E_{n-1}}{\rm{max}(E_{n},1)}\right|,$ (5)
$\Delta S_{n}$ $=\left|\frac{S_{n}-S_{n-1}}{S_{n}}\right|.$ (6)
Here $E_{n}$ and $S_{n}$ are the ground state energy and entanglement entropy at the middle of the lattice at sweep $n$. The denominator in $\Delta E_{n}$ is chosen to avoid numerical overflow if $E_{n}\approx 0$, which is not necessary in the case of entropy. In our program, we set maximum energy and entropy errors of $10^{-5}$ and $10^{-3}$, respectively, then when the wave function has errors above these bounds the optimization stops, which showed to be enough to achieve convergence.